# Ambiguity Clustering: an accurate and efficient decoder for qLDPC codes

### v1

@@ 3 Gaussian elimination and BP-OSD
The state of the art for qLDPC decoding is BP with Ordered Statistics Decoding post-processing (BP-OSD). BP-OSD uses the output of BP to inform an application of Gaussian elimination, a standard procedure in linear algebra that transforms a matrix $H$ into a form particularly convenient for solving equations of the form $He=\sigma$.
@@
A column of a matrix is in reduced form if it contains exactly one $1$. An $m\times n$ matrix as a whole is in reduced form if there is a set of $m$ columns in reduced form which have their $1$s in distinct rows.
A matrix in reduced form differs from the special form $H=\begin{bmatrix}I&B\end{bmatrix}$ only by row and column permutations, so is equally convenient for parameterising solutions to $He=\sigma$. This will be a repeating theme: a special form of $H$ will be presented with a particularly convenient ordering of the rows and columns, but any reordering of the special form will work equally well.
@@ 3.1 Gaussian elimination
The basic unit of Gaussian elimination is the pivot operation (Figure 4). Choose a pivot row $i$ and pivot column $j$ of $H$ that are adjacent (meaning $H_{ij}=1$) and add the $i$th row to every other row $\ell$ adjacent to column $j$ (with $H_{\ell j}=1$). After this pivot operation, the $j$th column is in reduced form. Moreover, since the rest of the column is now 00, further pivot operations will not take the $j$th column out of reduced form, as long as we avoid pivoting again in the same row.
The description of Gaussian elimination is now very short: for as long as we can choose a valid pivot entry $ij$ (with $H_{ij}=1$, and neither row $i$ nor column $j$ having previously been used as a pivot), pivot at $ij$. After $m$ pivot operations, $H$ will be in reduced form.
We can now decode as follows. Let $H$ be a parity check matrix, and let $H^{\prime}$ be a reduced form of $H$ obtained via Gaussian elimination. Since $H^{\prime}$ is obtained from $H$ by performing row operations, there is some invertible matrix $R$ such that $H^{\prime}=RH$. Given a syndrome $\sigma$, let $\sigma^{\prime}=R\sigma$. Then
@@
so we have an equivalent decoding problem to the one we started with, but where the parity check matrix is in reduced form. We can now explore the solution space to $H^{\prime}e=\sigma^{\prime}$ using the technique of (5). From now on, whenever we perform a pivot operation on $H$, we perform the same set of row operations on $\sigma$ to preserve the solution space to $He=\sigma$ (see Figure 4). We shall also omit the primes from $H^{\prime}$ and $\sigma^{\prime}$, using $H$ and $\sigma$ to refer to the current transformed state of the problem at any point.
@@ 3.2 Ordered Statistics Decoding
It generally remains impractical to iterate over all $2^{n-m}$ solutions to $He=\sigma$. A search would be much easier if the likely values of $e$ were almost entirely supported on the pivot columns. Ordered Statistics Decoding (OSD) aims to achieve this by exploiting our freedom to choose a pivot during Gaussian elimination: in each round, we select a pivot $ij$ which maximises $p_{j}$, BP’s estimate of the probability that $e_{j}=1$ given that $He=\sigma$. This is typically described as reordering the columns so that $p_{j}$ decreases left to right, then choosing the leftmost available pivot at each stage.
BP-OSD then is a family of algorithms of the following general form.
@@
Both the accuracy and run time can vary depending on the exact implementation of each stage. We will give a rough indication of the relative cost of each stage.
BP consists of some number of rounds $r$ passing messages along the edges of the Tanner graph. The Tanner graph has $m+n$ vertices. For a qLDPC decoding problem each vertex has bounded degree, so there are $O(m+n)=O(n)$ edges, and BP runs in time at most $O(rn)$. BP is highly amenable to parallelisation, so it is reasonable to think that this could be reduced to $O(r)$ in principle.
Gaussian elimination performs at most $m$ pivot operations, each consisting of at most $m-1$ additions of length $n$ rows, for a total complexity of $O(m^{2}n)$. The matrices involved in qLDPC decoding are sparse, so the practical complexity could be lower even if we make no special effort to use the sparsity, for example if each pivot operation involves relatively few rows.
The cost of the search stage depends heavily on how thoroughly we search. There are several named strategies in the literature [15].
@@
  • Combination sweep order $t$ (BP-OSD-CS($t$)). Try all $g$ of weight at most $1$, and those $g$ with exactly two $1$s in positions corresponding to the $t$ most likely non-pivot columns.
@@
Often $t$ is taken to be a small constant, giving satisfactory performance on small codes. However, the existence of trapping sets and the split belief problem (see Section 2) gives a heuristic argument that the amount of searching required by BP-OSD should grow exponentially in the size of the problem. If there are $s$ problematic areas of the syndrome, each of which can be resolved in at least two different ways, then an unstructured search must examine at least $2^{s}$ solutions to find the best one. But if these configurations occur at some constant rate, then for large problems $s$ will scale with $n$ for some constant fraction of syndromes supplied to the decoder.
@@ 4 Ambiguity Clustering
We now describe AC in full. Broadly speaking, AC works by performing incomplete Gaussian elimination on $H$, with modified pivot selection rules and modified stopping criteria. This has two advantages: first, it is cheaper than full Gaussian elimination. Second, the incomplete Gaussian elimination imposes a block structure on $H$, which allows for a more efficient search over the solution space.
After running BP to obtain estimated posteriors $p_{j}$, AC proceeds in three stages. In the first stage (Section 4.1), we obtain an initial solution to $He=\sigma$. In the second stage (Section 4.2), we grow outwards from this initial solution to identify additional errors, not present in the initial solution, that could be relevant to solving the decoding problem. Considering these additional errors builds a block structure in $H$, or equivalently a cluster structure in the Tanner graph. In the third stage (Section 4.3), we analyse each of the clusters separately.
We provide two figures to help the reader orient themselves. In Figure 5, we illustrate the evolution of the problem through stages 1 and  2. In Figure 7, we give a high-level overview of the algorithm.
@@
Figure 5: a. A general, unstructured, parity check matrix $H$. The syndrome $\sigma$ is shown as a column vector on its right.  b. The associated Tanner graph. Checks with a $1$ in the syndrome are marked in red.  c. After stage 1, $H$ is reduced with respect to the syndrome (Definition 4.1): an identity block covers the rows in which the syndrome has a $1$. In general, the rows and columns of the identity ‘block’ may be scattered throughout the matrix. Because a pivot row can still be involved in later pivot operations, there might (as in this example) be rows in the identity block without a $1$ in the syndrome.  d. In the Tanner graph, being reduced with respect to $\sigma$ means that every marked check has an associated error node adjacent only to that check.  e. After stage 2, an expanded set of blocks covers the syndrome. The blocks $C_{i}$ are shown as coloured rectangles. Each consists of an $m_{i}\times m_{i}$ identity matrix $I_{i}$, and optionally some linearly dependent columns in the $m_{i}\times(n_{i}-m_{i})$ matrix $B_{i}$. Again, in practice these rows and columns may be scattered throughout the matrix.  f. In the Tanner graph, each cluster consists of at least one check node–error node pair, indicated with an arrow, and optionally some additional errors adjacent only to those checks ‘gluing them together’ into a cluster. Every marked check is part of some cluster.
@@ 4.1 AC Stage 1: initial solution
The parity check matrix $H$ is reduced with respect to $\sigma$ if every row $i$ of the matrix with $\sigma_{i}=1$ is a pivot row.
If $H$ is reduced with respect to $\sigma$, then we have (7) up to permutations of the rows and columns. In comparison to the fully reduced form obtained in Section 3, we have done less work to find the first solution to $He=\sigma$. In return, we no longer have easy access to further solutions. Recovering these will be the purpose of stage 2.
To reduce $H$ with respect to $\sigma$, we choose pivots such that each pivot operation makes a row $i$ with $\sigma_{i}=1$ into a pivot row (possibly at the cost of creating additional $1$s in the syndrome). If we have multiple options, we break ties using the $p_{j}$ as in OSD.
@@
      the row must have a non-zero syndrome ($\sigma_{i}=1$); and
@@
    Perform a pivot operation using pivot row $i$ and pivot column $j$, updating both $H$ and $\sigma$ to preserve the solution space of $He=\sigma$.
@@
The key difference from OSD is in condition 1(c). In OSD, the order in which pivots are chosen is fixed in advance. By contrast, here we select the next pivot dynamically based on both $p_{j}$ and the current state of the updated syndrome. Since the original syndrome is sparse, this condition and the associated early-stopping behaviour allow the initial solution to be obtained much more quickly than in BP-OSD.
@@ 4.2 AC stage 2: cluster formation
In stage 2 we transform the matrix $H$ into its final form, a block structure that captures the local structure of the solution space. The blocks can also be viewed as clusters in a Tanner graph.
The block form we are aiming for is, up to row and column permutations,
@@
An example matrix of this form is shown in Figure 5. If we reinterpret $C$ as the adjacency matrix of a Tanner graph, the $c$ blocks correspond to connected components. Figure 5 shows these connected components embedded inside the full Tanner graph represented by $H$.
@@
with each $\sigma_{i}$ of length $m_{i}$. If we let
@@
for each $i$. The condition (13) is reasonable provided the zero block corresponds to error mechanisms that are unlikely to have occurred. As long as we can ensure this, we will have split the decoding problem into many smaller decoding problems that can be solved separately.
The partially reduced form (7) can be viewed as a trivial example of this block structure with $C$ consisting of the pivot rows and columns, with $m_{i}=n_{i}=1$ throughout. In stage 2 we will add an additional $K$ columns to $C$, and typically also some rows. Increasing $K$ increases the accuracy of the decoder at the cost of increased run time.
@@
Grow existing blocks. Otherwise, column $j$ is supported entirely on the rows of $C$, so we can add it to $C$ unchanged. It is possible that column $j$ has $1$s in several different blocks of $C$. In that case, we merge those blocks into a single larger block. The number of rows is unchanged, the number of columns increases by 1, and the number of blocks either stays the same or decreases.
@@
    Add column $j$ to $C$, either by seeding a new block or growing existing blocks as appropriate.
@@
    Repeat steps 1 and 2 until a total of $K$ additional columns have been added to $C$.
@@ 4.3 AC stage 3: cluster analysis
where we have labelled by $L_{i}$ the part of $L$ restricted to the columns in block $C_{i}$, and are promising to make no further use of the starred columns.
The triples $C_{i},\sigma_{i},L_{i}$ express smaller decoding problems which we shall solve separately. It turns out that many of these problems are very easy.
@@ 4.3.1 Unambiguous clusters
$\sum_{e_{i}\in\{0,1\}^{n_{i}}\colon C_{i}e_{i}=\sigma,\,L_{i}e_{i}=\lambda_{i}}\epsilon_{i}(e_{i}).$ (16)
@@
It turns out that there is frequently only a single value of $\lambda$ for which (16) is non-zero. That is, there is only a single logical effect consistent with the local syndrome, once we have committed to explaining that part of the syndrome locally. This is analogous to the neutral clusters arising during execution of the Union-Find decoder [6]: all corrections within a cluster are logically equivalent, so we can choose between them arbitrarily (for example, by ‘peeling’ in Union-Find).
@@ 4.3.2 Ambiguous clusters
In many cases, all clusters arising from stage 2 are unambiguous and no further analysis is necessary. When there is at least one ambiguous cluster, we have typically observed exactly one.
@@
As in Section 3.1, we let $e_{i}=\begin{bmatrix}f_{i}\\ g_{i}\end{bmatrix}$, with $f_{i}$ of length $m_{i}$ and $g_{i}$ of length $n_{i}-m_{i}$ so that the $2^{n_{i}-m_{i}}$ solutions to
@@ 4.3.3 Combining cluster results
4.4 Summary of stages of AC
Figure 7: Schematic overview of AC.
To summarise, AC has four stages, depicted in Figure 7:
  • Run BP on the syndrome to obtain posterior probability estimates $p_{j}$.
  • AC stage 1, initial solve. Informed by the syndrome and the BP posteriors, perform pivot operations to put the matrix $H$ into reduced form, sufficient to identify one solution.
  • AC stage 2, cluster formation. Perform further pivot operations and include linearly dependent errors to form a block structure.
  • AC stage 3, cluster analysis. For each logically ambiguous cluster, analyse a subset of the solutions to determine its likely logical effect, and sum these together with the effects of the unambiguous clusters to obtain a final overall logical effect.
@@ 5 Numerical methods
Bravyi et al. [14] described circuits to perform syndrome extraction for their codes. We use versions of these circuits made available by Gong et al. [23] for Gidney’s Clifford simulator Stim [26]. The circuits describe the preparations, gates and measurements for a quantum memory experiment: initialise a block of qubits, extract $r$ rounds of measurement data, then measure the physical qubits to determine the error states of the $k$ logical $Z$ observables.
We use the same circuit-level noise model as [14] with a single parameter $p$.
@@
  • After single qubit gates. With probability $p$, apply Pauli $X$, $Y$ or $Z$ with equal probability. (This is equivalent to depolarising the qubit with probability $34p/$ 3.) An idle qubit in any time step experiences a noisy identity gate.
@@
  • After two qubit gates. With probability $p$, apply one of the $15$ non-trivial 2-qubit Pauli operations $IX$, $IY$, $IZ$, $XI$, etc. (This is equivalent to depolarising both qubits with probability $1516p/$ 15.)
Stim performs two functions. First, it analyses the circuit to produce a ‘detector error model’, which contains information equivalent to the parity check matrix $H$, logical matrix $L$ and prior probabilities $\epsilon$. Second, it efficiently simulates the noisy syndrome extraction circuit, producing syndrome data and associated true logical effects against which we can assess the accuracy of a decoder.
The codes we test against encode more than one logical qubit. We count as a failure any run of the decoder in which we do not compute the full logical state correctly. We report the per-round logical failure rate
$p_{\text{fail}}=\frac{f}{rN}$ (20)
@@
where $N$ is the number of calls to the decoder, $f$ is the number of decoding failures, and $r$ is the number of rounds of syndrome extraction (in all cases set equal to the reported distance of the code). This formula ignores the second order effects of cancelling logical errors. These effects are more complicated to analyse for codes encoding multiple logical qubits than for those with a single logical qubit, but also less likely. The per-round logical failure rate for BP-OSD was computed in the same way from data kindly provided by the authors of [14].
@@
Timing data for both AC and BP-OSD were obtained by running each single-threaded on an M2-based MacBook Pro. We used a common Python harness that generates syndrome data using Stim then calls out to either our C++ implementation of AC, or the standard BP-OSD implementation by Roffe [15, 16] in Cython. Only time spent in the decoder is included in the reported per-round timing data.
Both AC and BP-OSD are families of decoding algorithms with various adjustable parameters.
In AC, we set

### v2

@@ 2 Gaussian elimination and BP-OSD
The state of the art for qLDPC decoding is BP with Ordered Statistics Decoding post-processing (BP-OSD). BP-OSD uses the output of BP to inform an application of Gaussian elimination, a standard linear algebra procedure that transforms a matrix $H$ into a form convenient for solving equations of the form $He=\sigma$.
@@
A column of a matrix is in reduced form if it contains exactly one $1$. An $m\times n$ matrix is in reduced form if there is a set of $m$ columns in reduced form with their $1$s in distinct rows.
A matrix in reduced form differs from the form $H=\begin{bmatrix}I&B\end{bmatrix}$ only by row and column permutations, so is equally convenient for solving $He=\sigma$. This will be a repeating theme: a special form of $H$ will be presented with a particularly convenient ordering of the rows and columns, but any reordering will work equally well.
@@ 2.1 Gaussian elimination
The basic unit of Gaussian elimination is the pivot operation (Figure 4). Choose a pivot row $i$ and pivot column $j$ of $H$ that are adjacent (in the corresponding Tanner graph, meaning $H_{ij}=1$) and add the $i$th row to every other row $\ell$ adjacent to column $j$ (with $H_{\ell j}=1$). After this pivot operation, the $j$th column is in reduced form. Moreover, since the rest of the column is now 00, further pivot operations will not take the $j$th column out of reduced form, as long as we avoid pivoting again on the same row.
The description of Gaussian elimination is now very short: for as long as we can choose a valid pivot entry $ij$ ($H_{ij}=1$, neither row $i$ nor column $j$ has previously been a pivot), pivot at $ij$. After $m$ pivot operations, $H$ is in reduced form.
We can now decode as follows. Let $H^{\prime}$ be a reduced form of $H$ obtained via Gaussian elimination. Since $H^{\prime}$ is obtained from $H$ by performing row operations, there is some invertible matrix $R$ recording those row operations such that $H^{\prime}=RH$. Given a syndrome $\sigma$, let $\sigma^{\prime}=R\sigma$. Then
@@
so we have an equivalent decoding problem to the one we started with, but with the parity check matrix in reduced form. We can now explore the solution space to $H^{\prime}e=\sigma^{\prime}$ using the technique of (5). From now on, whenever we perform a pivot operation on $H$, we perform the same set of row operations on $\sigma$ to preserve the solution space to $He=\sigma$ (see Figure 4). We also omit the primes from $H^{\prime}$ and $\sigma^{\prime}$, using $H$ and $\sigma$ to refer to the current transformed state of the problem at any point.
@@ 2.2 Ordered Statistics Decoding
It generally remains impractical to iterate over all $2^{n-m}$ solutions to $He=\sigma$. Searching would be easier if the likely values of $e$ were almost entirely supported on the pivot columns. Ordered Statistics Decoding (OSD) aims to achieve this by exploiting our freedom to choose a pivot during Gaussian elimination: in each round, we select a pivot $ij$ which maximises $p_{j}$, BP’s estimate of the probability that $e_{j}=1$ given that $He=\sigma$. This is typically described as reordering the columns so that $p_{j}$ decreases left to right, then choosing the leftmost available pivot at each stage.
BP-OSD then is a family of algorithms of the following form.
@@
The accuracy and run time can vary depending on the exact implementation of each stage. We will give a rough indication of the relative cost of each stage.
BP consists of some number of rounds $r$ passing messages along the edges of the Tanner graph. The Tanner graph has $m+n$ vertices. For qLDPC decoding problems each vertex has bounded degree, so there are $O(m+n)=O(n)$ edges, and BP runs in time at most $O(rn)$. BP is highly amenable to parallelisation, so it is reasonable to think that this could be reduced to $O(r)$ in principle.
Gaussian elimination performs at most $m$ pivot operations, each consisting of at most $m-1$ additions of length $n$ rows, for a total complexity of $O(m^{2}n)$. Matrices for qLDPC decoding are sparse, so the practical complexity could be lower even if we make no special effort to use the sparsity, for example if each pivot operation involves relatively few rows.
The cost of searching depends heavily on how thoroughly we search. There are several named strategies in the literature [49].
@@
  • Combination sweep order $t$ (BP-OSD-CS($t$)). Try all $g$ of weight at most $1$, and those $g$ with exactly two $1$s in the $t$ most likely non-pivot columns.
@@
Often $t$ is taken to be a small constant, giving satisfactory performance on small codes. However, the existence of trapping sets and the split belief problem gives a heuristic argument that the amount of searching required by BP-OSD should grow exponentially in the size of the problem. If there are $s$ problematic areas of the syndrome, each of which can be resolved in at least two different ways, then an unstructured search must examine at least $2^{s}$ solutions. But if these configurations occur at some constant rate, then for large problems $s$ will scale with $n$ for some constant fraction of syndromes.
@@ 3 Ambiguity Clustering
Here we detail the three stages of AC: stage 1 (Section 3.1), where we find an initial solution; stage 2 (Section 3.2), where we form complex clusters; and stage 3 (Section 3.3), where we analyse these clusters independently.
We provide two figures to help the reader orient themselves. In Figure 5, we illustrate the evolution of the problem through stages 1 and  2. Figure 7 contains a high-level overview of the algorithm.
@@
Figure 5: Evolution of the decoding problem through AC stages 1 and 2.a. A general, unstructured, parity check matrix $H$. The syndrome $\sigma$ is shown as a column vector on its right.  b. The associated Tanner graph. Checks with a $1$ in the syndrome are marked in red.  c. After stage 1, $H$ is reduced with respect to the syndrome (Definition 3.1): an identity block covers the rows in which the syndrome has a $1$. In general, the rows and columns of the identity ‘block’ may be scattered throughout the matrix. Because a pivot row can still be involved in later pivot operations, there might (as in this example) be rows in the identity block without a $1$ in the syndrome.  d. In the Tanner graph, being reduced with respect to $\sigma$ means that every marked check has an associated error node adjacent only to that check.  e. After stage 2, an expanded set of blocks covers the syndrome. The blocks $C_{i}$ are shown as coloured rectangles. Each consists of an $m_{i}\times m_{i}$ identity matrix $I_{i}$, and optionally some linearly dependent columns in the $m_{i}\times(n_{i}-m_{i})$ matrix $B_{i}$. Again, in practice these rows and columns may be scattered throughout the matrix.  f. In the Tanner graph, each cluster consists of at least one check node–error node pair, indicated with an arrow, and optionally some additional errors adjacent only to those checks ‘gluing them together’ into a cluster. Every marked check is part of some cluster.
@@ 3.1 AC Stage 1: initial solution
Having performed some sequence of pivot operations on distinct rows, the parity check matrix $H$ is reduced with respect to $\sigma$ if every row $i$ of the matrix with $\sigma_{i}=1$ is a pivot row.
If $H$ is reduced with respect to $\sigma$, then we have (7) up to permutations of the rows and columns. In comparison to the fully reduced form of Section 2, we have done less work to find the first solution to $He=\sigma$. In return, we no longer have easy access to further solutions. Recovering these will be the purpose of stage 2.
To reduce $H$ with respect to $\sigma$, we choose pivots such that each pivot operation makes a row $i$ with $\sigma_{i}=1$ into a pivot row (possibly at the cost of creating additional $1$s in the syndrome). We break ties using the $p_{j}$ as in OSD.
@@
      the row must have a non-zero syndrome ($\sigma_{i}=1$);
@@
    Perform a pivot operation using pivot row $i$ and pivot column $j$, updating both $H$ and $\sigma$ to preserve the solutions to $He=\sigma$.
@@
The key difference from OSD is condition 1(c). In OSD, the order in which pivots are chosen is fixed in advance. By contrast, here we select the next pivot dynamically based on both $p_{j}$ and the current state of the updated syndrome. Since the original syndrome is sparse, this allows the initial solution to be obtained much more quickly than in BP-OSD.
@@ 3.2 AC stage 2: cluster formation
In stage 2 we transform the matrix $H$ into its final form, a block structure that captures the local structure of the solution space. The blocks can also be viewed as clusters in the associated Tanner graph.
The form we are aiming for is, up to row and column permutations,
@@
An example matrix of this form is shown in Figure 5. If we reinterpret $C$ as the adjacency matrix of a Tanner graph, the $c$ blocks correspond to connected components. Figure 5 shows these components embedded inside the full Tanner graph represented by $H$.
@@
with each $\sigma_{i}$ of length $m_{i}$. If we restrict attention to $e$ of the form
@@
for each $i$. The restriction (13) asserts that columns outside $C$ corresponds to error mechanisms which did not occur. If this is not reliably satisfied, then AC will lose access to solutions that could have been found by OSD. If it is satisfied, we have split the decoding problem into many smaller decoding problems that can be solved separately.
The partially reduced form (7) is a trivial example of this block structure with $C$ consisting of the pivot rows and columns, with $m_{i}=n_{i}=1$ throughout. In stage 2 we will add an additional $K$ columns to $C$, and typically also some rows. Increasing $K$ increases the accuracy of the decoder at the cost of increased run time.
@@
Grow existing blocks. Otherwise, column $j$ is supported entirely on the rows of $C$, so we can add it to $C$ unchanged. It is possible that column $j$ has $1$s in multiple blocks of $C$. In that case, we merge those blocks into a single larger block. The number of rows is unchanged, the number of columns increases by 1, and the number of blocks either stays the same or decreases.
@@
    Add column $j$ to $C$, seeding a new block or growing existing blocks as appropriate.
@@
    Repeat steps 1 and 2 until $K$ additional columns have been added to $C$.
@@ 3.3 AC stage 3: cluster analysis
where we have labelled by $L_{i}$ the part of $L$ restricted to the columns in block $C_{i}$, and promise to make no further use of the starred columns.
The triples $C_{i},\sigma_{i},L_{i}$ express smaller decoding problems which we shall solve separately. It turns out that many of these problems are easy.
@@ 3.3.1 Unambiguous clusters
$\sum_{e_{i}\in\{0,1\}^{n_{i}}\colon C_{i}e_{i}=\sigma_{i},\,L_{i}e_{i}=\lambda_{i}}\epsilon_{i}(e_{i}).$ (16)
@@
It turns out that there is frequently only one value of $\lambda_{i}$ for which (16) is non-zero. That is, there is a unique logical effect consistent with the local syndrome, once we have committed to explaining that part of the syndrome locally. This is analogous to the neutral clusters arising during execution of the Union-Find decoder [6]: all corrections within a cluster are logically equivalent, so we can choose between them arbitrarily (for example, by ‘peeling’ in Union-Find).
@@ 3.3.2 Ambiguous clusters
In many cases, all clusters arising from stage 2 are unambiguous and no further analysis is necessary. An ambiguous cluster must be ‘large’ in the sense that it can be resolved in at least two different ways differing by a logical operation. We might expect that it would be difficult to arrange two large but disconnected clusters, and when running the decoder we typically see at most one ambiguous cluster.
@@
As in Section 2.1, we let $e_{i}=\begin{bmatrix}f_{i}\\ g_{i}\end{bmatrix}$, with $f_{i}$ of length $m_{i}$ and $g_{i}$ of length $n_{i}-m_{i}$, so that the $2^{n_{i}-m_{i}}$ solutions to
@@ 4 Numerical methods
Bravyi et al. [9] described circuits to perform syndrome extraction for their codes. We use versions of these circuits made available by Gong et al. [46] for Gidney’s Clifford simulator Stim [54]. The circuits describe a quantum memory experiment: initialise a block of qubits, extract $r$ rounds of measurement data, then measure the qubits to determine the states of the $k$ logical $Z$ observables. The decoding problem corresponding to a circuit is described in Section A.3. As in [9], we simulate the full syndrome extraction circuit, but only use $Z$-stabiliser measurements to decode the $Z$-logicals.
We use the circuit-level noise model from [9].
@@
  • After single qubit gates. With probability $p$, apply Pauli $X$, $Y$ or $Z$ with equal probability. An idle qubit in any time step experiences a noisy identity gate.
@@
  • After two qubit gates. With probability $p$, apply one of the $15$ non-trivial 2-qubit Pauli operations $IX$, $IY$, $IZ$, $XI$, etc. with equal probability.
We report the per-round logical failure rate
$p_{\text{fail}}=\frac{N_{\text{fails}}}{rN_{\text{shots}}}$ (20)
@@
where $N_{\text{shots}}$ is the number of syndromes decoded, $N_{\text{fails}}$ is the number of times we did not recover the full logical state correctly, and $r$ is the number of rounds of syndrome extraction (equal to the reported distance of the code). Second order effects of cancelling logical errors are more complicated to analyse than for codes with a single logical qubit, but also less likely. We compute $p_{\text{fail}}$ for BP-OSD in the same way from data kindly provided by the authors of [9].
@@
Timing data for both AC and BP-OSD were obtained by running each single-threaded on an M2-based MacBook Pro. A common Python harness generated syndrome data using Stim then called out to either our C++ implementation of AC, or the BP-OSD implementation by Roffe [49, 50] in Cython. Only time spent in the decoder is included in the per-round timing data.
We set the parameters of AC as follows. For each of 9, 19 and 29 rounds of (sum-product) BP we set $K$ (the number of additional columns added to $C$ in stage 2) equal to $\kappa n$, where $n$ is the number of columns of the parity check matrix, and increased $\kappa$ in steps of $0.01$ from 00 until AC matched or exceeded the accuracy of BP-OSD at each data point plotted in Figure 3. This gave three timings with comparable accuracy. The lowest timing was obtained for 9 rounds in almost all cases, so we used 9 rounds of BP throughout. Where we were able to decode at lower error rates than the authors of [9], so that there was no BP-OSD comparison data, we maintained the value of $\kappa$ selected for the lowest value of $p$ for which comparison data was available.
The timing data for BP-OSD was obtained using 10 000 rounds of (min-sum) BP and combination sweep order $t=7$, as in [9].
If BP alone was able to explain the syndrome, then the second stage decoder (OSD or AC) was not used.
In a previous version of this paper we set the number of rounds of BP for both BP-OSD and AC equal to the distance of the code. This was a bad choice that increased the run times of both BP-OSD and AC.
For BP-OSD, increasing the number of rounds of BP from $d$ to 10 000 decreases the overall running time by around a factor of 5, since it less often has to resort to the expensive OSD step.
Since AC is much faster than OSD, significantly increasing the number of rounds of BP was not advantageous for AC. Instead, we observed that sum-product BP exhibits strong periodic behaviour, such that the accuracy of AC is significantly increased if we use an odd number of rounds of BP. We could then trade off this increased accuracy to obtain a (less dramatic) reduction in run time.
Appendix A Setting up decoding problems
In this section we describe how the parity check matrix $H$, logical matrix $L$ and vector of prior probabilities $\epsilon$ of Section 1 can be obtained in various settings.
A.1 Classical binary linear codes
Let $H$ be the $(n-k)\times n$ parity check matrix of a classical binary linear code, and let
$G=\begin{bmatrix}I&\star\end{bmatrix}$ (21)
be a $k\times n$ generator matrix in standard form. The logical value of a codeword can be read out by projecting onto the first $k$ coordinates, so the readout of the $i$th encoded bit is affected precisely by errors in the $i$th message bit and we should take
$L=\begin{bmatrix}I&0\end{bmatrix}$ (22)
as the $k\times n$ logical matrix.
Typically the noise model is that each bit flips independently with probability $p$, so $\epsilon$ will be a constant vector with every entry equal to $p$.
A.2 Stochastic Pauli noise and perfect stabiliser measurements
Let $g_{1},\ldots,g_{m}$ be a choice of stabiliser generators for an $n$-qubit stabiliser code and let $q_{1},\ldots,q_{k}$ be a list of logical Pauli operators whose values we wish to decode. Let $a_{1},\ldots,a_{3n}=X_{1},Y_{1},Z_{1},X_{2},\ldots,Z_{n}$ be an enumeration of the non-identity single qubit Pauli errors. Then the $m\times 3n$ parity check matrix $H$ is defined by
$H_{ij}=\begin{cases}1&\text{if }a_{j}\text{ anticommutes with }g_{i},\\ 0&\text{otherwise.}\end{cases}$ (23)
Similarly, the $k\times 3n$ logical matrix $L$ is defined by
$L_{ij}=\begin{cases}1&\text{if }a_{j}\text{ anticommutes with }q_{i},\\ 0&\text{otherwise.}\end{cases}$ (24)
Typically the noise model is thought of as each qubit depolarising independently with some fixed probability $p$. As described in Section B, this is equivalent to each single qubit Pauli error occurring independently with some other probability $q$, so again $\epsilon$ is typically some vector with constant entries.
A.3 Circuit-level noise
A Clifford circuit is a circuit consisting of qubit preparations in the $X$- or $Z$-bases, Clifford gates, and measurements of Pauli operators. Clifford circuits are efficiently simulable using the stabiliser formalism, for example using Gidney’s Stim [54].
Running (or simulating) a Clifford circuit produces a list of measurement results, which can be identified with a vector in $\mathbb{F}_{2}^{N}$ for some $N$. Not all measurement sequences in $\mathbb{F}_{2}^{N}$ are achievable: for example, repeated measurement of the same stabiliser will produce the same outcome. In fact, the achievable measurement sequences form an affine subspace $V$ of $\mathbb{F}_{2}^{N}$. If we run a Clifford circuit with noise, deviation of the measurement sequence from $V$ is evidence of an error having occurred.
In the context of a quantum memory experiment, the linear equations defining this subspace some in two types. A detector is a set of measurement results whose sum takes a known value under error-free conditions, and which we are allowed to use to diagnose errors. A logical observable is also a set of measurement results whose sum takes a known value under error-free conditions, except that this knowledge relies on us having set up the experiment to preserve a logical |$\ket{0}$, say, in which case a measurement of the logical $Z$ operator would take a known value. Logical observables cannot be used for error correction, as in a practical application of quantum error-correction the logical state would be unknown.
For concreteness, we describe the situation for a small quantum memory experiment for a distance 3 rotated surface code with 3 rounds of syndrome extraction. We might simulate the following Clifford circuit.